# Semi-supervised and unsupervised learning for health indicator extraction from guided waves in aerospace composite structures

### v1

@@ 4 Discussion
By analysing Diversity-DeepSAD and DTC-VAE, key differences in their performances and behaviors are revealed. These differences shed light on their respective strengths and limitations in extracting meaningful HIs for composite structure prognostics.
@@
Diversity-DeepSAD graphs show a tendency for HIs to increase in sharp steps, which could indicate different damage states during the fatigue life of composite structures. However, these increases were sometimes erroneous or abrupt; this is most visible in Figure 8 in test Sample 3 at $50\text{\,}\mathrm{kHz}$ and test Sample 4 at $125\text{\,}\mathrm{kHz}$, showing the model may be sensitive to sudden changes in test sample structural health or sensor damage. On the other hand, DTC-VAE did not demonstrate the same behaviour, showing a smoother increase over the specimen lifetime, indicating improved performance in this regard. This is particularly the case with FFT features, with some exceptions in the case of HT features.
@@
Furthermore, results from both DTC-VAE and in particular Diversity-DeepSAD showed a tendency for HIs to spike at EoL, such as test Sample 3 at $125\text{\,}\mathrm{kHz}$. This is due to the model learning the failure state as a separate category to the continuous range of degradation up until that point, an effect which can be emphasised with alternative hyperparameters. This suggests that the deviation in the structural health condition after failure is significantly greater and distinct from the condition before failure, demonstrating the damage accumulation process in composite structures. Although high prognosability scores may indicate that extracted HIs are more suitable for RUL prediction, it is crucial to ensure that changes in HI values are not too abrupt. For example, a sharp increase near the end of life may leave insufficient time to take maintenance action. In this regard, DTC-VAE could potentially provide HIs that are more useful than Diversity-DeepSAD.
@@
The incremental changes in HIs produced by Diversity-DeepSAD could on the one hand be seen as a more interpretable result, potentially illustrating the distinct damage states of composite structures. This could provide valuable insights for prognostic models, especially state-based ones, in predicting RUL. However, the increases in HIs should not be too abrupt, and the overall trend must meet the criteria; otherwise, using HIs for RUL prediction would be challenging, as discussed previously. Further studies and testing are needed to establish a stronger connection between these phases and actual physical damage states. On the other hand, DTC-VAE resulted in overall smoother HIs, with slightly higher fitness scores and less sharp jumps, that could improve RUL prediction.
@@
Based on the average prognostic criteria and fitness scores across all units, Diversity-DeepSAD and DTC-VAE demonstrate competitive performance compared to other proposed models, as shown in Table 11. In this table, two SHM techniques—guided wave (GW) and acoustic emission (AE)—are also compared. It should be noted that the constructed HIs using AE are inherently time-dependent due to the nature of the sensing technology, whereas the GW technique has the potential to provide data for designing historical-independent HIs, as demonstrated by the proposed models in this study and HT-SSCNN [15].
@@
While SSLSTM using AE data [23] achieves the highest fitness score across all units ($F_{all}$) with a performance of 93%, it relies on test units for validation during the LOOCV process to stop training and fine-tune hyperparameters. This reliance potentially limits its applicability to real-world scenarios where test units are new and unknown. However, this limitation has been addressed in subsequent work [24], where CEEMDAN-SSLSTM followed by a BiLSTM ensemble achieved the best performance for AE data, with 91.3% for $F_{all}$ and 86.3% for test units ($F_{test}$).
@@
On the other hand, using GW data, the proposed historical-independent HIs in this study result in the lowest deviation between fitness scores obtained for all units ($F_{all}$) and test units ($F_{test}$), demonstrating reduced overfitting. Both the semi-supervised Diversity-DeepSAD model and the unsupervised DTC-VAE model are significantly more stable, with much lower standard deviations compared to existing state-of-the-art models. The DTC-VAE model, followed by WAE fusion, achieved the highest performance of 92% for test units ($F_{test}$) and 92.3% for all units ($F_{all}$). Notably, the deviation between fitness scores for all units and test units is minimal (0.3%).
@@
Table 11: Performance of various methods in the ReMAP dataset collected using guided wave (GW) and acoustic emission (AE) techniques.
@@
HT-SSCNN [15]
SSLSTM [23]
@@
CEEMDAN-SSLSTM [24]
@@
SHM technique GW AE
@@ 5 Conclusion
This study proposed a novel plenary data-driven framework for constructing health indicators (HIs) for aeronautical composite structures by developing two advanced machine learning models: Diversity-DeepSAD and DTC-VAE. First, statistical features from the best-performing signal processing techniques, fast Fourier transform (FFT) and Hilbert transform (HT), were extracted as inputs for the AI models. This step enabled dimensionality reduction of guided wave signals, each 2000 data points in length, recorded by PZTs across 56 actuator-sensor inspection paths at six excitation frequencies (2000×56×6 data points per measurement), to just a few tens of features per excitation frequency. This allows for simpler deep learning architectures while preserving critical information. Feature selection was based on fitness scores, calculated as the sum of the prognostic criteria of monotonicity, prognosability, and trendability. FFT features demonstrated superior performance, achieving a performance of 71.0%, i.e., an average fitness score of 2.13 (out of 3.00), compared to 68.2% for HT features.
@@
Diversity-DeepSAD and DTC-VAE exhibited distinct behaviors in HI construction. Diversity-DeepSAD produced incremental changes in HIs, potentially reflecting distinct damage states, but occasionally showed abrupt increases, particularly near the end of life (EoL). In contrast, DTC-VAE generated smoother HIs with higher fitness scores, benefiting from its monotonicity-constrained loss function, which directly optimized the fitness metric. This smoother trend makes DTC-VAE more suitable for remaining useful life (RUL) prediction, as it avoids abrupt changes that could limit actionable maintenance time. Therefore, the augmented Diversity-DeepSAD is recommended as a closer representative of true health state, while DTC-VAE may be more useful for prognostic applications.
@@
The proposed framework outperformed state-of-the-art models in constructing historical-independent HIs. DTC-VAE, followed by weighted averaging ensemble fusion, achieved the highest test fitness score of 2.76 (92% performance) with minimal deviation (0.3%) from the fitness score across all units, indicating reduced overfitting and improved generalization. Moreover, the models demonstrate greater stability in performance compared to existing approaches. Unlike some existing models, which rely on test units for validation, the proposed framework ensures applicability to real-world scenarios where test units are unknown.
Overall, this study provides an efficient and robust framework for HI construction, leveraging the strengths of Diversity-DeepSAD and DTC-VAE. These models enhance the accuracy and stability of HIs while supporting more reliable RUL predictions, contributing to improved maintenance strategies and operational safety for aeronautical composite structures. Future work will focus on extending the DTC-VAE loss function to incorporate additional prognostic criteria and validating the framework on diverse datasets to further enhance its applicability.
@@ Appendix A List of features by domain
Table 12: Time domain features [23].
@@
Table 13: Frequency domain features [23].
@@
Table 14: Time-frequency domain features [66].

### v2

@@ 5 Discussion
This section contextualises the results by analysing the behaviour of the constructed HIs, assessing their robustness and computational feasibility, and comparing the proposed models with state-of-the-art methods. The discussion links model behaviour to underlying physical mechanisms and highlights practical implications for GW SHM.
5.1 Interpretation of constructed HIs
Developing robust and reliable HIs for composite structures remains a challenging task, as these indicators must capture meaningful trends while remaining interpretable and suitable for prognostic applications. This analysis evaluates the augmented Diversity-DeepSAD and DTC-VAE, highlighting key differences in their performances and behaviours, which reveal their respective strengths and limitations in extracting meaningful HIs for composite structure prognostics.
@@ 5.1 Interpretation of constructed HIs
Diversity-DeepSAD graphs show a tendency for HIs to increase in sharp steps, which could indicate different damage states during the fatigue life of composite structures. However, these increases were sometimes erroneous or abrupt; this is most visible in Figure 10 in test Sample 3 at $50\text{\,}\mathrm{kHz}$ and test Sample 4 at $125\text{\,}\mathrm{kHz}$, showing the model may be sensitive to sudden changes in test sample structural health or sensor damage. On the other hand, DTC-VAE produced smoother HIs, indicating improved performance in this regard. This is particularly the case with FFT features, with some exceptions in the case of HT features.
The incremental changes in HIs produced by Diversity-DeepSAD could be seen as a more explainable result, potentially illustrating the distinct damage states of composite structures. This could provide valuable insights for prognostic models, especially state-based ones, in predicting RUL. However, the increases in HIs should not be too abrupt, else using HIs for RUL prediction would be challenging, as discussed previously. Further studies and testing are needed to establish a stronger connection between these phases and actual physical damage states. On the other hand, HIs produced by DTC-VAE resulted in slightly higher fitness scores and less sharp jumps, that could improve RUL prediction.
Figure 15: Post-buckling stiffness against number of compression cycles for each specimen.
In order for further physical interpretation, Figure 15 shows the stiffness of each specimen, which is estimated by the slope of the linear region of the load–displacement curves during post-buckling (i.e. ranging from 40 kN to 60 kN) [Yue2022]. The stiffness reduction is due to complex damage mechanisms and therefore also represents the physical health of the structure, similarly to the generated HIs. It should be noted that stiffness was not included in the data fed to the models, as it cannot be easily measured in operation, particularly in an aerospace application.
The HIs produced by augmented Diversity-DeepSAD and DTC-VAE can be meaningfully linked to plausible damage mechanisms and measured stiffness trends, while remaining cautious about causality (see Table 11). Stepwise HIs produced by Diversity-DeepSAD highlight discrete changes in GW scattering—consistent with localised events such as delamination initiation, disbond propagation, or sudden coalescence of cracks—whereas the DTC-VAE yields smoother HIs that reflect the gradual accumulation of distributed microdamage (e.g. matrix cracking). Considering the impact/disbond locations, that of Sample 2 lies off the stiffener and shows the earliest HI transition but the smoothest stiffness decline, suggesting local damage detected early by GW but not yet affecting global stiffness; Sample 4’s is located on the stiffener and exhibits long stability followed by late HI jumps and a sharper stiffness loss, consistent with critical stiffener-related propagation; Sample 5’s pre-existing disbond explains its early, abrupt stiffness reduction ($\sim$20% lifetime) while GW responses are path-dependent and thus more model-sensitive. Where both models and feature folds consistently indicate transitions (for instance Sample 2’s early change), confidence is increased that the HI reflects a genuine physical change. Differences in timing between HI transitions and stiffness drops therefore provide diagnostic value: early HI changes with late stiffness loss point to local damage not yet structural, while early stiffness drops with variable GW response point to gross load-path alterations (e.g. a disbond).
Table 11: Summary of data-driven HI behaviour (constructed by augmented Diversity-DeepSAD vs. DTC-VAE, both upon FFT features), stiffness evolution, impact/disbond location, and plausible physical interpretation.
Specimen
Impact/disbond
& relative position
DeepSAD’s HI
(damage states before
final jump to EoL & timing)
DTC-VAE’s HI
(shape & timing)
Stiffness behaviour Plausible physical interpretation (linking HI → stiffness)
Sample 1
Impact near
stiffener edge
2 damage states;
first state longer than second
Very smooth HI;
gradual monotonic rise
Smoothed sudden
reduction, later in
life
Impact near stiffener edge causes early local cracking/small
delaminations detected by GW as two discrete stages (DeepSAD).
Global stiffness declines later when local damage coalesces.
DTC-VAE’s smooth HI captures overall monotonic accumulation.
Sample 2
Impact at
off-stiffener region
3 damage states;
earliest first transition
among samples;
long final state
Fast HI rise early
(first $\sim$30% lifetime),
then slow growth
Smoothest stiffness
decline; no early
jump (stiffness
accelerates near
EoL)
Off-stiffener impact produces local scattering changes that GW
detects very early (DeepSAD). Because damage is local and not
initially in primary load paths, global stiffness remains unaffected
until late accumulation—consistent with DTC-VAE plateau after
early rise.
Sample 3
Impact near
opposite stiffener edge
4 damage states;
3rd state longest
Gradual multi-stage
trend, intermediate
between extremes
Smooth reduction;
late acceleration
toward EoL
Multiple HI states suggest sequential damage mechanisms (matrix
cracking → interface debonding → delamination) at/around the
stiffener edge. Stiffness reduces progressively and accelerates as
damage becomes global.
Sample 4
Impact on
the stiffener, center
2 damage states;
fewer pre-final stages
Slow progression until
$\sim$70% lifetime, then jump(s)
toward EoL
Early-to-mid life
smoother drop
then sharper loss
Impact on stiffener produces damage that remains subcritical for
long time (little HI change) then propagates rapidly along the stiffener
causing HI jumps and stiffness loss (loss of load-transfer at stiffener).
Timing matches DTC-VAE late jumps.
Sample 5
Manufacturing disbond
between stiffener and
skin (right side)
2 damage states (sometimes
3 depending on fold);
early transition visible
Smooth monotonic HI
but early rise relative
to some units
Earliest stiffness
decrease
($\sim$20% lifetime);
sudden early drop
Pre-existing disbond alters global load path → early measurable
stiffness loss. GW HIs detect local scattering but effectiveness depends
on which paths intersect the disbond; model variability reflects path
sensitivity. Overall, early stiffness drop is consistent with a structural
discontinuity (disbond) rather than purely fatigue accumulation.
@@
Furthermore, results from both DTC-VAE and in particular Diversity-DeepSAD showed a tendency for HIs to spike at EoL, such as test Sample 3 at $125\text{\,}\mathrm{kHz}$. This is due to the model learning the failure state as a separate category to the continuous range of degradation until that point, an effect which can be emphasised with alternative hyperparameters. This suggests that the deviation in the structural health condition after failure is significantly greater and distinct from the condition before failure, demonstrating the damage accumulation process in composite structures. Although high prognosability scores may indicate that extracted HIs are more suitable for RUL prediction, it is crucial to ensure that changes in HI values are not too abrupt. For example, a sharp increase near the end of life may leave insufficient time to take maintenance action. In this regard, DTC-VAE could potentially provide HIs that are more useful than Diversity-DeepSAD.
@@
5.2 Uncertainty and dataset size
The suitability of incorporating uncertainty quantification into the evaluation of the constructed HIs was examined, given its common application for prognostic modelling. Although it is widely used in RUL prediction [Wang2024UC], its use in HI construction remains limited. Studies on HIs typically rely on the use of prognostic criteria and do not report confidence intervals [Moradi2024, Yang2023]. This is partly because HIs represent abstract latent variables rather than directly measurable physical quantities. As a result, interval-based uncertainty estimates for fitness scores may be misleading, especially in relation to smaller datasets where statistical variance cannot be reliably defined. Nevertheless, as reported in this study, evaluating the standard deviation of $Mo$, PrPr, TrTr, and the fitness score—across (i) multiple random initialisations of the deep learning models, which reflects model uncertainty, and (ii) different composite specimens, which reflects variability due to GW sensing, manufacturing imperfections, and minor differences in sensor placement and test setup—is necessary. These statistics provide a quantitative assessment of the robustness of the learned HIs and their sensitivity to both model and experimental variability.
Although only five composite specimens were employed in this study, the dataset is highly information-rich: each unit provides dozens of GW measurements, each comprising 56 actuator–sensor paths across six frequencies. Since the proposed models operate in a history-independent manner, every GW measurement timestep forms an independent sample, yielding approximately 230 machine learning measurements in total. Furthermore, the leave-one-unit-out evaluation, where an entire specimen is unseen during training, provides a stringent generalisation test and helps mitigate overfitting despite the limited number of structural units.
5.3 Computational feasibility
Both Diversity-DeepSAD and DTC-VAE remain computationally lightweight. All experiments were conducted on a laptop equipped with a 12th-Gen Intel Core i7-12700H CPU (14 cores, 20 threads), 16 GB RAM and an NVIDIA RTX A1000 Laptop GPU with 4 GB VRAM. Training and inference were performed using GPU acceleration (CUDA 12.8). Training a single model instance, in other words one frequency and feature type for one validation fold, required on average $35.68\text{\,}\mathrm{s}$ for Diversity-DeepSAD and $2.27\text{\,}\mathrm{s}$ for DTC-VAE, with a maximum of $62.34\text{\,}\mathrm{s}$ and $3.01\text{\,}\mathrm{s}$ respectively. The corresponding parameter counts averaged 519,664 trainable parameters for Diversity-DeepSAD and 71,309 for DTC-VAE (1.982 MB and 0.272 MB respectively in memory), depending mainly on the input dimensionality and hidden layer width.
Inference for a new specimen, i.e. generating a complete HI trajectory for a single specimen at one frequency, took on average $17.47\text{\,}\mathrm{ms}$ per forward pass for Diversity DeepSAD and 1.03 ms for DTC-VAE. Since training and Bayesian optimisation are performed offline, only this inference step is relevant for online SHM. The sub-second inference times leave a large margin with respect to typical GW measurement intervals, indicating that both methods are suitable for real-time deployment, for example, for in-service aircraft structures in the industrial context. Moreover, the training time scales approximately linearly with the number of specimens and epochs, and model size scales linearly with the hidden layer width, suggesting that the framework can be extended to larger datasets and additional frequencies without prohibitive computational overhead.
5.4 Comparison with state-of-the-art
@@
Based on the average prognostic criteria and fitness scores across all units, Diversity-DeepSAD and DTC-VAE demonstrate competitive performance compared to other proposed models, as shown in Table 12. In this table, two SHM techniques—GW and acoustic emission—are also compared. It should be noted that the constructed HIs using acoustic emission are inherently time-dependent due to the nature of the sensing technology, whereas the GW technique has the potential to provide data for designing history-independent HIs, as demonstrated by the proposed models in this study and HT-SSCNN [Moradi2024].
@@
While SSLSTM using acoustic emission data [Moradi2023Intelligent] achieves the highest fitness score across all units ($F_{all}$) with a performance of 93%, it relies on test units for validation during the LOOCV process to stop training and fine-tune hyperparameters. This reliance potentially limits its applicability to real-world scenarios where test units are new and unknown. However, this limitation has been addressed in subsequent work [moradi2025novel], where CEEMDAN-SSLSTM followed by a BiLSTM ensemble achieved the best performance for acoustic emission data, with 91.3% for $F_{all}$ and 86.3% for test units ($F_{test}$).
@@
On the other hand, using GW data, the proposed history-independent HIs in this study result in the lowest deviation between fitness scores obtained for all units ($F_{all}$) and test units ($F_{test}$), demonstrating reduced overfitting. Both the semi-supervised Diversity-DeepSAD model and the unsupervised DTC-VAE model are significantly more stable, with much lower standard deviations compared to existing state-of-the-art models. The DTC-VAE model, followed by WAE fusion, achieved the highest performance of 92% for test units ($F_{test}$) and 92.3% for all units ($F_{all}$). Notably, the deviation between fitness scores for all units and test units is minimal (0.3%).
@@
Table 12: Performance of various methods in the ReMAP dataset collected using guided wave (GW) and acoustic emission techniques.
@@
HT-SSCNN [Moradi2024]
SSLSTM [Moradi2023Intelligent]
@@
CEEMDAN-SSLSTM [moradi2025novel]
@@
SHM technique GW Acoustic Emission
Although this study focuses on the model families previously explored for the ReMAP GW dataset, the results highlight opportunities for evaluating more advanced architectures. Models such as transformers, graph neural networks, and physics-informed neural networks may further exploit spatial–temporal dependencies in GW signals. A systematic assessment of these architectures is left for future work.
5.5 Limitations and future work
Although the proposed framework demonstrates strong performance and robustness, some limitations identified in the discussion are summarised explicitly:
  1. (i)
    The dataset contains only five composite specimens. The data richness and leave-one-unit-out evaluation provide sufficient validation coverage, however, future work should aim to include more specimens to improve statistical generalisation.
  2. (ii)
    Diversity-DeepSAD and DTC-VAE occasionally produce abrupt transitions at EoL, somewhat influencing the prognosability criterion and their use for prognostics. Further work should aim to incorporate further constraints into the learning process and loss functions.
  3. (iii)
    The framework relies on multi-frequency guided wave measurements with dense actuator–sensor paths, which may not be available in all SHM deployments.
  4. (iv)
    Only two model families were explored. More advanced architectures, such as transformers, graph neural networks, and physics-informed neural networks, may further improve degradation representation.
@@ 6 Conclusion
This study proposed a novel plenary data-driven framework for constructing health indicators (HIs) for aeronautical composite structures by developing two advanced machine learning models: Diversity-DeepSAD and DTC-VAE. First, statistical features from the best-performing signal processing techniques, fast Fourier transform (FFT) and Hilbert transform (HT), were extracted as inputs for the AI models. This step enabled dimensionality reduction of GW signals, each 2000 data points in length, recorded by PZTs across 56 actuator-sensor inspection paths at six excitation frequencies (2000×56×6 data points per measurement), to just a few tens of features per excitation frequency. This allows for simpler deep learning architectures while preserving critical information. Feature selection was based on fitness scores, calculated as the sum of the prognostic criteria of monotonicity, prognosability, and trendability. FFT features demonstrated superior performance, achieving a performance of 71.0%, i.e. an average fitness score of 2.13 (out of 3.00), compared to 68.2% for HT features.
@@
Diversity-DeepSAD and DTC-VAE exhibited distinct behaviours in HI construction. Diversity-DeepSAD produced incremental changes in HIs, potentially reflecting distinct damage states, but occasionally showed abrupt increases, particularly near the end of life (EoL). In contrast, DTC-VAE generated smoother HIs with higher fitness scores, benefiting from its monotonicity-constrained loss function, which directly optimised the fitness metric. This smoother trend makes DTC-VAE more suitable for RUL prediction, as it avoids abrupt changes that could limit actionable maintenance time. Therefore, the augmented Diversity-DeepSAD is recommended as a closer representative of true health state, while DTC-VAE may be more useful for prognostic applications.
@@
The proposed framework outperformed state-of-the-art models in constructing history-independent HIs. DTC-VAE, followed by weighted averaging ensemble fusion, achieved the highest test fitness score of 2.76 (92% performance) with minimal deviation (0.3%) from the fitness score across all units, indicating reduced overfitting and improved generalisation. Moreover, the models demonstrate greater stability in performance compared to existing approaches. Unlike some existing models, which rely on test units for validation, the proposed framework ensures applicability to real-world scenarios where test units are unknown.
Overall, this study addresses the critical challenge of constructing robust HIs for aeronautical composite structures, where reliable monitoring and prognostics are essential for operational safety and effective maintenance planning. By leveraging the strengths of augmented Diversity-DeepSAD and DTC-VAE, the proposed frameworks provide stable HIs while supporting more reliable RUL predictions. The frameworks are designed to be computationally efficient, scalable and feasible for monitoring, making them suitable for SHM applications such as in-service aircraft structures. Future work will focus on extending the DTC-VAE loss function to incorporate additional prognostic criteria, exploring larger datasets with more units to validate generalisability and assessing the implementation in operational environments to further enhance the frameworks practical applicability. Additional work will also explore the applicability of graph neural networks and physics-informed neural networks to GW SHM.
@@ Appendix A List of features by domain
Each time or time-frequency domain feature $X$ is given in terms of the time or time-frequency domain signal samples $x$, while frequency domain features $S$ are given in terms of the frequency domain signal samples $s(f)$,
Table 13: Time domain features [Moradi2023Intelligent].
@@
Table 14: Frequency domain features $S$ [Moradi2023Intelligent].
@@
Table 15: Time-frequency domain features $X$ [Buckley2023].
@@ 4 Kurtosis $X_{kurtosis}=\frac{\Sigma^{N}_{i=1}\left(x(i)-X_{m}\right)^{4}}{(N-1)X_{sd}^{4}}$
Appendix B Sensitivity Analysis
All plots for the sensitivity analysis described in Subsection 3.7 are included below in Figure 16 and Figure 17.
(a) Diversity-DeepSAD using FFT features.
(b) Diversity-DeepSAD using HT features.
Figure 16: FUll hyperparameter sensitivity analysis of Diversity-DeepSAD over $(\nu,\eta,\lambda)$, showing response surfaces for the two SP methods.
(a) DTC-VAE using FFT features.
(b) DTC-VAE using HT features.
Figure 17: Full hyperparameter sensitivity analysis of DTC-VAE over $(\alpha,\beta,\gamma)$, showing response surfaces for the two SP methods.
Appendix C Code Availability
The source code used for all experiments in this study is available on GitHub at the following repository:
https://github.com/mortezamkh1992/Diversity-DeepSAD-vs-DTC-VAE